Session: Oct 2005 (session 3)
Wavelength: 21 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Monari at Thu Nov 3 18:40:35 2005) Minor failures :
Some scans lost due to errors in positioning in schedule

Onsala (Michael Lindqvist at Thu Nov 3 19:03:10 2005) Minor failures : WIND
Antenna stowed 07:30-10:30 due to high winds. Bad weather, rain, high winds (tracking problems). Source at very low elevation at the end ==> high tsys.

Torun (Kaz Borkowski at Fri Nov 4 06:39:58 2005) Success :
No known problems except that towards the end a source went below the telescope horizon.

Westerbork (Tony Foley at Fri Nov 4 10:32:34 2005) Minor failures : SCHED
Wb used in phased array mode. Some scans at the end out of hourangle range. We also lost 11:48-12:12UT when try to restart after a gap when we replaced some array IF hardware giving problems.

Torun (Kaz Borkowski at Mon Nov 7 09:27:10 2005) Failure : RECO
The entire record has been inadvertedly erased (with sstest program) while trying to cope with problems in recording on the same 8-pack of a later experiment (see feedback at GM057B).

Arecibo (T.Ghosh/E.Momjian/C.J.Salter at Tue Nov 8 14:04:45 2005) Minor failures :
BBC#8 with lower-sideband level-setting problem (seen during the GL028 runs) was physically interchanged with the BBC#5. Since the OH-Mainline data was streamed through to the location of the BBC#5, and as far as we know, this shource didn't have much signal in the mainlines, this seemed to us to be a better solution. In addition, we lowered the levels of the IF signals to all BBCs, and monitored the BBC#5 -- (containig the faulty unit at that point) during the entire run -- levels seemed stabler. We think that there is a very good chance that data on all bands will correlate for this run.

Effelsberg (A. Kraus at Wed Nov 9 16:39:41 2005) Success :

Jodrell_Bank (Alastair Gunn at Thu Nov 10 12:30:41 2005) Success :
The Lovell Telescope was used. No known problems.